# Spatial-Mamba: Effective Visual State Space Models via Structure-aware State Fusion

### v1

@@ Appendix D More Qualitative Results
In this section, we present the visualization results for object detection and instance segmentation tasks in Fig. 7 and for semantic segmentation task in Fig. 8. Compared with VMamba, our Spatial-Mamba demonstrates superior performance in both tasks, producing more accurate detections with higher confidence levels and more refined segmentations, particularly in areas where local structure information is crucial. For example, in the second row of Fig. 7, VMamba mistakenly identifies the shoes on a skateboard as a person, likely because observing the shoes from only four directions makes them resemble a human. Our method avoids this mistake by simultaneously perceiving the shoes and their surrounding context. Similarly, in the semantic segmentation task, as shown in the second and third rows of Fig. 8, our approach achieves more precise segmentations of the structure of trees and doors. These results highlight the effectiveness of our proposed Spatial-Mamba in leveraging local structural information for better visual understanding.
@@
Figure 7: Visualization examples for object detection and instance segmentation tasks on COCO dataset with Mask R-CNN 1$\times$ schedule between VMamba and Spatial-Mamba.
@@
Figure 8: Visualization examples for semantic segmentation task on ADE20K dataset with single-scale inputs between VMamba and Spatial-Mamba.

### v2

@@ Appendix E Qualitative Results
In this section, we present the visualization results of object detection and instance segmentation in Fig. 7, and present the results of semantic segmentation in Fig. 8. Compared with VMamba, our Spatial-Mamba demonstrates superior performance in both tasks, producing more accurate detection boxes segmentation masks, particularly in areas where local structural information is crucial. For example, in the second row of Fig. 7, VMamba mistakenly identifies the shoes on a skateboard as a person, probably because it observes the shoes from four directions independently and resembles them as a human. Our method avoids this mistake by simultaneously perceiving the shoes and their surrounding context. Similarly, in the semantic segmentation task, as shown in the second and third rows of Fig. 8, our approach achieves more precise structures of trees and doors. These results highlight the effectiveness of our proposed Spatial-Mamba in leveraging local structural information for better visual understanding.
@@
Figure 7: Visualization examples of object detection and instance segmentation on COCO dataset with Mask R-CNN 1$\times$ schedule (He et al., 2017) by VMamba and Spatial-Mamba.
@@
Figure 8: Visualization examples of semantic segmentation on ADE20K dataset with single-scale inputs by VMamba and Spatial-Mamba.
Appendix F Further Discussions about SASF
SASF extension. To further explore and validate the spatial modeling capability of our proposed SASF, we adapt SASF into Vim and VMamba backbones, resulting in Vim+SASF (by replacing the middle class token with average pooling and setting state dimension to 1) and VMamba+SASF (directly integrated with SASF). Under the same training settings, the classification results on ImageNet-1K are presented in Tab. 7. It can be seen that by integrating with SASF, the bi-directional Vim model is improved by 0.5% in top-1 accuracy and the cross-scan VMamba is improved by 0.3%. These findings verify that SASF can even enhance the performance of multi-directional models, which have already incorporated (yet in an inefficient manner) spatial modeling operations.
Table 7: Comparison of classification performance on ImageNet-1K by integrating with SASF, where ‘Throughput’ is measured using an A100 GPU with an input resolution of $224\times 224$.
[…]
Vim-Ti $224^{2}$ 6M 1.4G - 71.7
Vim-Ti+SASF $224^{2}$ 7M 1.6G - 72.2
VMamba-T $224^{2}$ 30M 4.9G 1686 82.6
VMamba-T+SASF $224^{2}$ 31M 5.1G 1126 82.9
Fusion operators. There are also different choices of the operators that can be used for fusing state variables in SASF. We opt for depth-wise convolution due to its simplicity and computational efficiency. However, operators like dynamic convolution (Chen et al., 2020), deformable convolution (Dai et al., 2017), and attention mechanisms have also demonstrated superior performance in computer vision tasks. Therefore, we anticipate that they can be used as alternatives to the depth-wise convolution in our Spatial-Mamba. We simply train a Spatial-Mamba-T network with dynamic convolution and the results are shown Tab. 8. We can see that dynamic convolution yields a 0.2% performance gain in accuracy but significantly reduces throughput from 1438 to 907. This result suggests the potential advantages of more flexible fusion operators for SASF, but also highlights the importance of considering computational cost.
Table 8: Comparison of classification performance by Spatial-Mamba-T with depth-wise conv. and dynamic conv. on ImageNet-1K, where ‘Throughput’ is measured using an A100 GPU with an input resolution of $224\times 224$.
[…]
Depth-wise Conv. $224^{2}$ 27M 4.5G 1438 83.5
Dynamic Conv. $224^{2}$ 33M 4.8G 907 83.7
Appendix G Effective Receptive Field (ERF)
We compare the Effective Receptive Field (ERF) (Ding et al., 2022) of the center pixel on popular backbone networks before and after training, as shown in Fig. 9. The ERF values represent the contributions of every pixel on input space to the central pixel in the final output feature maps. To visualization, we randomly select 50 images from the ImageNet-1K validation set, resize them to a resolution of 1024$\times$1024, and then calculate the ERF values with the auto-grad mechanism. Before training, our Spatial-Mamba-T initially exhibits a larger receptive field than other methods except DeiT-S due to neighborhood connectivity in the state space. After training, our method, along with DeiT-S, Vim-S, and VMamba-T, all demonstrate a global ERF. In addition, both Vim-S and VMamba-T exhibit noticeable accumulation contributions along either horizontal or vertical directions, which can be attributed to their multi-directional fusion mechanisms. In contrast, our unidirectional Spatial-Mamba-T effectively eliminates this directional bias.